# A Systematic Literature Review on Detecting Software Vulnerabilities with Large Language Models

### v1

@@ 4.1.2. Additional Objectives
The developments in additional objectives reflect a broader shift in the role of LLMs from simple classifiers to end-to-end vulnerability assistants that support developers interactively throughout the software development life cycle. The presented objectives also create opportunities for multitask benchmarks that evaluate detection accuracy alongside, e.g., reasoning quality, localization precision, and severity estimation, offering a more holistic perspective on vulnerability remediation.
4.2. Input Representation
In the context of LLM-based vulnerability detection, the way source code and related information are represented as input plays a crucial role in model performance. Beyond plain software code, inputs may include natural language vulnerability descriptions or structured abstractions of code. These representations influence the model’s ability to capture both syntax and deeper semantic relationships relevant to vulnerability patterns. Building on the taxonomy introduced for vulnerability repair by Zhang et al. (Zhang et al., 2024a), we adapt it for the vulnerability detection task, categorizing input representations as raw, structure-aware, prompt, and conversation-style.
Raw
Raw input representations of code involve directly feeding source code into the model, without any additional annotations or transformations (Ahmad et al., 2021; Atiiq et al., 2024b, a; Buratti et al., 2020; Cao et al., 2024; Chan et al., 2023; Chen et al., 2024, 2023; Cipollone et al., 2025; Curto et al., 2024a, b; d’Aloisio et al., 2025; Deng et al., 2024; Ding et al., 2025; Du et al., 2023; Ehrenberg et al., 2024; Ferrag et al., 2025b; Fu et al., 2024; Fu and Tantithamthavorn, 2022; Fu et al., 2023a; Gao et al., 2023b; Gonçalves et al., 2025; Grishina et al., 2023; Prabith et al., 2024; Gujar, 2024; Guo et al., 2024b; Hanif and Maffeis, 2022; Ghafoor Hussain et al., 2025; Islam et al., 2024b, a; Jiang et al., 2024b, 2025, c; Ji et al., 2024; Kalouptsoglou et al., 2024a, b; Kong et al., 2024; Le et al., 2025, 2024; Li et al., 2025f; Lin et al., 2022a; Li et al., 2025b; Liu et al., 2024e; Lu et al., 2023; Luo et al., 2021, 2025; Mamede et al., 2022; Manuel et al., 2024; Mao et al., 2025; Mechri et al., 2025; Mei et al., 2023; Nguyen et al., 2024; Pan et al., 2021; Qi et al., 2024; Rahman et al., 2024; Ridoy et al., 2024; Risse and B”̈ohme, 2024; Russo et al., 2025; Safdar et al., 2023; Shestov et al., 2025; Shimmi et al., 2024a; Singh et al., 2022; Steenhoek et al., 2025b, 2023b; Sultana et al., 2024; Sultan et al., 2024; Su and Wu, 2023; Tian et al., 2024; Torkamani et al., 2025; Uddin et al., 2022; Wang et al., 2025c, 2024b, 2024c, 2024d, 2023e; Wen et al., 2024b; Xiong and Dong, 2024; Xu et al., 2023; Yang et al., 2025c, a, 2024a; Yin et al., 2024; Yuan et al., 2022; Yusuf and Jiang, 2024; Zeng et al., 2023; Zhang et al., 2025b, 2024d; Zhao et al., 2024a; Zhou et al., 2024a, b; Zhu et al., 2024; Ziems and Wu, 2021; Akli et al., 2025; Bappy et al., 2025; Hin et al., 2022; Ibanez-Lissen et al., 2025; Kaanan et al., 2024; Mahmud et al., 2025; Melo et al., 2025; Tian et al., 2025; Wen et al., 2023; Zheng et al., 2025; Lan et al., 2025; Panichella, 2025; Shi et al., 2024; Weyssow et al., 2025). This approach treats code as plain text sequences, relying entirely on the model’s pre-trained understanding of programming languages to detect patterns indicative of vulnerabilities. This representation is widely used for training or fine-tuning LLMs. However, raw code inputs may struggle to capture the deeper semantics and context necessary for detecting complex or subtle vulnerabilities. As such, this approach may underperform in scenarios requiring precise reasoning about program logic, data flow, or code structure.
Structure-Aware
Structure-aware representations enhance the model’s understanding of code by capturing deeper syntactic and semantic structures of code and vulnerabilities, thereby addressing the limitations of relying solely on textual patterns. Code is transformed into graph-based or other structured formats that expose control flow, data flow, and dependency relationships relevant to vulnerability semantics.
@@ 4.2. Input Representation
Common graph representations include Abstract Syntax Trees (ASTs) (Ni et al., 2023b; JianJie and Le, 2023; Sahar et al., 2024; Lu et al., 2024; Bahaa et al., 2024; Wen et al., 2024a; Wang et al., 2024e; Ahmed et al., 2025; Xue et al., 2025), Control Flow Graphs (CFGs) (Zhang et al., 2023a; Lu et al., 2024; Sun et al., 2024), Data Flow Graphs (DFGs) (Yu et al., 2022; Liu et al., 2024f; Li et al., 2025d), Program Dependency Graphs (PDGs) (Lu et al., 2024; Du et al., 2024; Chen et al., 2025), combining data and control dependencies, and Code Property Graphs (CPGs) (Liu et al., 2025c; Feng et al., 2025; Yang et al., 2025b; Rangapuram and Ratha, 2025), which integrate multiple code views (AST, CFG, PDG). Some studies employ composite or multi-layered graph structures that integrate several of the above (Wang et al., 2023b; Islam et al., 2023; Zheng et al., 2024; Liu et al., 2025a), which can outperform other graphs (Zhang et al., 2023b), or combine multiple modalities including text, graphs, and visual representations (Ni et al., 2023a).
@@
In addition to graph structures, other structured inputs aim to isolate semantically relevant code regions, such as program slices (Zhu et al., 2022; Ding et al., 2023; Huang et al., 2022; Sun et al., 2023; Lingxiang et al., 2025; Halder et al., 2025), e.g., meaningful blocks based on AST structure (Lingxiang et al., 2025), or code gadgets (Kim et al., 2022; Thapa et al., 2022; Omar, 2023; Omar and Shiaeles, 2023; G”̈urfidan, 2024; Li et al., 2021; Purba et al., 2023; Peng et al., 2023; Quan et al., 2025; Zhao and Guan, 2025). Code gadgets, in particular, consist of program statements that are semantically related to each other through data or control dependencies (Li et al., 2018), isolating vulnerability-relevant code regions. Other studies include program traces, i.e., execution-aware code structure (Ding et al., 2024b; Mächtle et al., 2025). Mächtle et al. (Mächtle et al., 2025) specifically build trace gadgets, i.e., trace the code before slicing. By making structural and semantic dependencies explicit, structure-aware representations enable models to reason more effectively about vulnerability-related behaviors, especially those that span across the codebase or require understanding of program logic and execution flow.
Prompt
Prompt engineering has emerged as a widely used technique for steering LLMs towards effective vulnerability detection without requiring any modifications to the model’s weights. Prompt engineering formulates the task through strategically designed input prompts, typically consisting of a system prompt (providing general context or behavioral instructions) and a user prompt (specifying the task and providing the code to be analyzed).
Prompt content often extends beyond a basic task instruction. It may include vulnerability reports or CWE descriptions (Cao and Jun, 2024; Du et al., 2025; Sheng et al., 2024; Shimmi et al., 2024b; Sun et al., 2025), function-level insights such as descriptions of program logic, control flow (Tamberg and Bahsi, 2025), or static analysis outputs (Keltek et al., 2024; Mohajer et al., 2024). Well-designed prompts are crucial to ensure the model receives relevant context and is guided toward the correct reasoning path.
Some studies engage the model in a back-and-forth analysis, rather than issuing a single prompt for vulnerability detection (Liu et al., 2024f; Gao et al., 2023a; Yu et al., 2024; Liu et al., 2023a; Ullah et al., 2024). This interactive, Conversation-Style setting enables the model to double-check and revise earlier predictions, aiming to mitigate as many vulnerabilities as possible through a process of progressive refinement (Liu et al., 2024c). Such settings support multi-step reasoning, for example, by discussing program analysis before initiating vulnerability detection (Li et al., 2024; Liu et al., 2023a; Espinha Gasiba et al., 2024; Wang et al., 2025b; Huynh et al., 2025). Conversation-style interactions are particularly prominent in self-refinement workflows, where models are prompted to review and improve their outputs (Yu et al., 2024; Du et al., 2024; Li et al., 2024; Tamberg and Bahsi, 2025; Eberhardt and Milánkovich, 2024).
A notable technique using prompt-based inputs (apart from typical prompt engineering) is prompt-tuning, which involves learning prompts to optimize model performance on a downstream task. Two main prompt forms are distinguished: hard prompts, which are manually crafted sequences of tokens appended to the input, and soft prompts, which are learned continuous embeddings that serve as virtual tokens prepended to the input.
4.3. System Architecture
@@
Vulnerability detection systems using LLMs differ significantly in how the models are integrated into the overall architecture. Broadly, we categorize existing approaches into two groups: LLM-centric systems, where the LLM serves as the core analytical component, and hybrid systems, where the LLM is used in conjunction with other models.
4.3.1. LLM-Centric
In LLM-centric approaches, the vulnerability detection task is handled primarily by the language model itself. These systems rely on the LLM for core tasks such as classification and reasoning, and typically use either prompt engineering or fine-tuning to adapt the model to the detection task. Surveyed studies rely on two main types of LLMs: general-purpose LLMs and code LLMs. Depending on the study setup, models from both categories may be used independently (comparison or benchmark) or in combination (ensemble), e.g., (Ridoy et al., 2024; Mahmud et al., 2025).
General-Purpose LLMs
General-purpose LLMs are pre-trained on natural language corpora and tend to perform well on explanation-based or natural language-heavy tasks. Early examples include BERT (Devlin et al., 2019) and its variants RoBERTa (Liu et al., 2019) and DistilBERT (Sanh et al., 2020). Other models widely used for vulnerability detection include: the GPT series (Radford et al., 2019; Brown et al., 2020), with the latest version being GPT-4 (OpenAI et al., 2024); the LLaMA series (Touvron et al., 2023a, b; Grattafiori et al., 2024); the Qwen series (Bai et al., 2023; Yang et al., 2024b; Qwen et al., 2025); Gemma 1 and 2 (Team et al., 2024b, c); DeepSeek R1 and v2 (DeepSeek-AI et al., 2025, 2024a); Mistral (Jiang et al., 2023) and Mixtral (Jiang et al., 2024a); T5 (Raffel et al., 2020); Claude 3 (Anthropic, 2024); Gemini 1.5 (Team et al., 2024a); and Phi2 (Javaheripi et al., 2023) and Phi3 (Abdin et al., 2024).
Code LLMs
Code LLMs are pre-trained specifically on large-scale code corpora for code-related tasks, enhancing performance on syntax-sensitive or program-structure-aware tasks.
Examples include: BERT-based models, such as CodeBERT (Feng et al., 2020) and GraphCodeBERT (Guo et al., 2021); StarCoder (Li et al., 2023), and StarCoder2 (Lozhkov et al., 2024); UniXcoder (Guo et al., 2022); the GPT-based Codex (Chen et al., 2021); CodeLlama (Rozière et al., 2024); CodeQwen 1.5 (Bai et al., 2023), Qwen2.5-Coder (Hui et al., 2024); CodeGemma (Team et al., 2024d); DeepSeek-Coder (Guo et al., 2024c; DeepSeek-AI et al., 2024b); and CodeT5(+) (Wang et al., 2021, 2023c).
4.3.2. Hybrid

### v2

@@ 4.1.2. Generation (F2)
These generative tasks reflect a broader shift in the role of LLMs from simple classifiers to end-to-end vulnerability assistants that support developers interactively throughout the software development life cycle. Figure 4 maps the relationships between classification and generation tasks across the surveyed studies. Each node is annotated with the absolute number of studies using that task formulation, providing a quantitative overview of the landscape. The flows further visualize studies that combine classification with generative tasks. These combined tasks create opportunities for multitask benchmarks that evaluate detection accuracy alongside, e.g., reasoning quality, localization precision, and severity estimation, offering a more holistic perspective on vulnerability remediation.
4.2. Input Representation (I)
The representation of the input code and additional context plays a pivotal role in effective vulnerability detection. Distinct from traditional static analysis tools that require language-specific rules, LLMs offer unique capabilities in how they process input. First, their extensive pre-trained world knowledge allows them to interpret diverse input forms, e.g., code and natural language instructions. This flexibility enables the processing of multi-language source code with varied vulnerability types and auxiliary information within a unified framework. Second, their semantic understanding allows LLMs to generalize; by reasoning about the input rather than matching fixed vulnerable patterns, LLMs may identify rare or novel vulnerabilities. Considering these capabilities, we categorize input representations along the code and auxiliary information axes.
4.2.1. Code (I1)
Raw (I1.1) input representations of code involve feeding source code directly into the model during inference as plain text sequences (used by 139 studies). This approach relies entirely on the model’s understanding and pre-learned knowledge of programming syntax to detect vulnerable patterns. This representation is widely used for training or fine-tuning LLMs. However, raw code inputs may struggle to capture the deeper semantics and context necessary for detecting subtle or complex vulnerabilities. As such, this approach may underperform in scenarios requiring precise reasoning about program logic, data flow, or code structure.
@@ 4.2.1. Code (I1)
Combining raw code with or expanding representations to Structured (I1.2) code enhances the model’s understanding of code by capturing deeper syntactic and semantic structures of code and vulnerabilities, thereby addressing the limitations of relying solely on textual patterns. Code is transformed into graph-based or other structured formats that expose control flow, data flow, and dependency relationships relevant to vulnerability semantics. Common graph representations include Abstract Syntax Trees (ASTs) (Ahmed et al., 2025; Bahaa et al., 2024; JianJie and Le, 2023; Li et al., 2025j; Wen et al., 2024a; Zhang et al., 2025a; Yu and Zhang, 2025), specifically flattened (Li et al., 2025j) or decomposed (Zhang et al., 2025a), Control Flow Graphs (CFGs) (Gajjar et al., 2025a; Liu et al., 2025a; Lu et al., 2024; Sun et al., 2024; Zhang et al., 2023a), Data Flow Graphs (DFGs) (Jiang et al., 2024b; Liu et al., 2024f; Wang et al., 2024b; Yu et al., 2022), forms of Program Dependency Graphs (PDGs) (Du et al., 2024; Hin et al., 2022; Lu et al., 2024; Sun et al., 2025c; Wu and Zhang, 2024), combining data and control dependencies, and Code Property Graphs (CPGs) (Chang et al., 2024; Feng et al., 2025; Lekssays et al., 2025; Li et al., 2025e; Liu et al., 2025b; Ngan et al., 2025; Ni et al., 2023a; Rangapuram and Ratha, 2025; Liang et al., 2024; Yu and Zhang, 2025), which integrate multiple code views (AST, CFG, PDG). Some studies employ composite or multi-layered graph structures that integrate several of the above (Jiang et al., 2022; Liu et al., 2024a; Islam et al., 2023, 2024a; Zheng et al., 2025, 2024; Liu et al., 2025a; Lu et al., 2024), which can outperform other graphs (Zhang et al., 2023b), or combine multiple modalities including text, graphs, and visual representations (Ni et al., 2023a).
@@
In addition to graph structures, other structured inputs aim to isolate semantically relevant code regions, such as program or code slices (Wu and Xiao, 2025; Tian and Zhang, 2025; Ding et al., 2023; Huang et al., 2022; Sun et al., 2023; Zhu et al., 2022), chunks (Halder et al., 2025), or code gadgets (Pham et al., 2024; Gujar, 2024; G”̈urfidan, 2024; Kim et al., 2022; Li et al., 2021; Omar, 2023; Omar and Shiaeles, 2023; Peng et al., 2023; Purba et al., 2023; Quan et al., 2025; Thapa et al., 2022; Zhao and Guan, 2025; Zhao et al., 2023; Chen et al., 2025a). Code gadgets, in particular, consist of program statements that are semantically related to each other through data or control dependencies (Li et al., 2018), isolating vulnerability-relevant code regions. Other studies include program traces, i.e., execution-aware code structure (Ding et al., 2024b; Mächtle et al., 2025). Mächtle et al. (Mächtle et al., 2025) specifically build trace gadgets, i.e., trace the code before slicing. By making structural and semantic dependencies explicit, structure-aware representations enable models to reason more effectively about vulnerability-related behaviors, especially those that span across the codebase or require understanding of program logic and execution flow.
4.2.2. Auxiliary Information (I2)
To enhance the effectiveness of vulnerability detection, studies move beyond source code by augmenting the input with auxiliary information, particularly in prompt-engineering settings. This additional information provides the model with domain knowledge or code context. We categorize this auxiliary information into (i) vulnerability information, (ii) semantic artifacts, (iii) execution artifacts, and (iv) tool output.
Vulnerability Information (I2.1) involves explicitly equipping the model with domain vulnerability knowledge. Most commonly, studies include a pre-defined list of CWEs to the prompt to narrow the classification scope (Bakhshandeh et al., 2023; Cao and Jun, 2024; Dozono et al., 2024; Fu et al., 2023b; Gnieciak and Szandala, 2025; Li et al., 2025i; Liu et al., 2024b; Sayagh and Ghafari, 2025; Shashwat et al., 2024; Steenhoek et al., 2025a; Wang et al., 2023a; Xia et al., 2025; Yu et al., 2025) or provide labeled examples (few-shot settings) (Chan et al., 2023; Ding et al., 2025; Dolcetti et al., 2025; Farr et al., 2025; Hannan et al., 2025; Li et al., 2025j; Liu et al., 2024f, 2023; Mao et al., 2025; Ni et al., 2025; Nong et al., 2024; Ouchebara and Dupont, 2025; Safdar et al., 2025a; Shu et al., 2025; Sovrano et al., 2025; Trad and Chehab, 2025; Tsai et al., 2025a; Wang et al., 2025b; Widyasari et al., 2024; Yang et al., 2025c; Yildiz et al., 2025; Yin et al., 2024; Zhou et al., 2024a, b; Chen et al., 2025a; Jiao et al., 2025; Tsai et al., 2025b; Le et al., 2024; Lu et al., 2024; Sheng et al., 2024; Zhang et al., 2025b). Other studies include more detailed vulnerability (CWE) definitions and descriptions (Sun et al., 2025b; Shimmi et al., 2024b; Li et al., 2025g; Kouliaridis et al., 2024; Ullah et al., 2024; Zhu et al., 2025a; Mohajer et al., 2024; Widyasari et al., 2025; Ceka et al., 2025; Pelofske et al., 2024; Mathews et al., 2024; Du et al., 2025a), particularly as part of retrieval architectures, or full analysis reports (Luo et al., 2025).
Semantic Artifacts (I2.2) focus on the high-level intent and functionality of the code, bridging the gap between syntax and understanding. For example, studies provide natural language explanations of the code functionality (Gajjar et al., 2025b; Torkamani et al., 2025), questions regarding the code’s behavior (Jararweh et al., 2025), explicit descriptions of data sources and sinks (Li et al., 2025e), or code comments and docstrings (Huynh et al., 2025).
Execution Artifacts (I2.3) capture the behavior, structural dependencies, or runtime environment of the code. Studies enrich inference input with data and control dependencies, interprocedural interactions, and environment constraints (Ahmed et al., 2025). For instance, researchers include information on callee functions (Sun et al., 2025b; Li et al., 2025i), dynamic execution signals such as unit test results (Mhatre et al., 2025), or code-context (Li et al., 2025f; Sayagh and Ghafari, 2025; Li et al., 2025i), e.g., global variables, type declarations, and library imports. Data flow analysis techniques are also used (Li et al., 2025j; Tamberg and Bahsi, 2025; Yang et al., 2025b; Zhou et al., 2025b), e.g., to extract dependencies from call graphs (Wen et al., 2024b).
Tool Output (I2.4) augments the model context with the outputs of other tools, most commonly SAST tools, allowing the LLM to interpret the static analysis findings to refine further predictions (Bakhshandeh et al., 2023; Keltek et al., 2024; Munson et al., 2025; Wang et al., 2025e; Yang et al., 2025c). Yang et al. (Yang et al., 2025c) inputs the results of other deep learning models (i.e., the prediction probability) to inform the final classification.
4.3. System Architecture (S)
@@
Vulnerability detection systems using LLMs differ significantly in which and how models are integrated into the overall system. Broadly, we categorize existing approaches into two groups: LLM-centric systems, where the LLM serves as the main predictive component, and hybrid systems, which integrate an LLM with other learning-based components.
4.3.1. LLM (S1)
In LLM-centric approaches, the LLM performs the core predictive tasks, such as classification and reasoning. The model is typically adapted to the detection task through prompt engineering or fine-tuning. Additional tools, such as SAST, only provide auxiliary information rather than predictive signals. The surveyed studies employ a variety of LLMs with different architectures, pre-training domains, and scales. In the following, we discuss LLMs within these categories to categorize systems in more detail.
Model Architecture (S1.1)
Most LLMs are based on the Transformer architecture, which includes encoder and decoder modules with a self-attention mechanism. We follow other works and categorize the LLMs used in the surveyed studies into the three architectural groups (i) encoder-only, (ii) encoder-decoder, and (iii) decoder-only, as introduced by Pan et al. (Pan et al., 2024).
Encoder-Only (S1.1.1) LLMs, e.g., models from the BERT (Devlin et al., 2019) family (CodeBERT (Feng et al., 2020), GraphCodeBERT (Guo et al., 2021), RoBERTa (Liu et al., 2019), DistilBERT (Sanh et al., 2020)), use only the encoder to encode the sentence and comprehend the relationships between words (Pan et al., 2024). The encoder maps the input code into a fixed-size dense vector representation (embedding) that captures semantic relationships. In vulnerability detection, these models are typically adapted via fine-tuning by adding a classification head (e.g., a multi-layer perceptron) to output a probability score or class label. Alternatively, they function as frozen feature extractors, generating embeddings that serve as input for external classifiers.
Encoder-Decoder (S1.1.2) LLMs, e.g., T5 (Raffel et al., 2020), CodeT5 (Wang et al., 2021, 2023c), or UniXcoder (Guo et al., 2022), use both the encoder and decoder module in a sequence-to-sequence framework (Pan et al., 2024). The encoder maps the input code into a latent hidden-state representation, which the decoder uses to generate a target output sequence. For vulnerability detection, this architecture allows the task to be framed as text generation, where the input is the source code and the output is a generated text string, e.g., the label ”vulnerable” or a natural language explanation.
Decoder-Only (S1.1.3) LLMs use only the decoder module, using auto-regressive mechanisms to predict the next token in a sequence. The input, typically a prompt containing instructions and the code snippet, is processed sequentially, and the output is generated as a continuation of this input, enabling the generation of reasoning chains and structured vulnerability reports. Many state-of-the-art LLMs follow this architecture, e.g., the GPT family (Radford et al., 2019; Brown et al., 2020; OpenAI et al., 2024) and the Llama family (Touvron et al., 2023a, b; Grattafiori et al., 2024). Decoder-only LLMs are the dominant architecture used for vulnerability detection in the surveyed studies.
Pre-Training Domain (S1.2)
Surveyed studies rely on LLMs from two main domains: general-purpose LLMs and code LLMs. General-Purpose LLMs (S1.2.1) are pre-trained on natural language corpora and tend to perform well on explanation-based or natural language-heavy tasks. Examples from this domain are the BERT family, the GPT family, the Llama family, the Qwen series (Bai et al., 2023; Yang et al., 2024b; Qwen et al., 2025), Gemma (Team et al., 2024b, c), DeepSeek (DeepSeek-AI et al., 2025, 2024a), Mistral (Jiang et al., 2023), Mixtral (Jiang et al., 2024a), Claude (Anthropic, 2024), Gemini (Team et al., 2024a), and Phi (Javaheripi et al., 2023; Abdin et al., 2024). Code LLMs (S1.2.2) are pre-trained specifically on large-scale code corpora for code-related tasks, enhancing performance on syntax-sensitive or program-structure-aware tasks. Examples include CodeBERT (Feng et al., 2020), GraphCodeBERT (Guo et al., 2021), StarCoder (Li et al., 2023; Lozhkov et al., 2024), UniXcoder (Guo et al., 2022), the GPT-based Codex (Chen et al., 2021), CodeLlama (Rozière et al., 2024), CodeQwen 1.5 (Bai et al., 2023), Qwen2.5-Coder (Hui et al., 2024), CodeGemma (Team et al., 2024d), DeepSeek-Coder (Guo et al., 2024c; DeepSeek-AI et al., 2024b), and CodeT5 (Wang et al., 2021, 2023c).
Model Scale (S1.3)
The parameter count of an LLM typically serves as a proxy for its capacity to learn complex patterns and encode domain knowledge. However, increasing scale imposes proportional demands on computational resources for training and inference. Notably, there exists no formal consensus regarding the exact parameter threshold that constitutes a ”large” language model. Despite this ambiguity, to analyze the trade-off between performance and resource efficiency, we categorize the models used in the surveyed studies into four distinct groups:
Tiny (S1.3.1): Models with fewer than 1 billion parameters (1B). This category primarily includes pre-trained language models such as CodeBERT or UniXcoder. Due to their lightweight nature, they are highly efficient for fine-tuning and inference, making them suitable for resource-constrained environments.
Small (S1.3.2): Models with 1 to 8 billion parameters (1B–8B), such as Gemma-2B, Llama2-7B, or Llama3.1-8B. This scale represents a trade-off for many researchers; these models provide a significant improvement in generative capabilities over tiny models while remaining small enough to be deployed or fine-tuned on high-end consumer-grade hardware.
Medium (S1.3.3): Models between 8 and 70 billion parameters (8B–70B), including CodeLlama-34B, Mixtral 8x7B, and DeepSeek-R1-32B. These models typically outperform smaller counterparts in complex reasoning tasks but exceed the memory capacity of standard consumer hardware, necessitating enterprise-grade GPUs or multi-GPU setups.
Large (S1.3.4): Models exceeding 70 billion parameters (70B), such as GPT-4, Llama3.1-405B, and DeepSeek-v3. These models achieve state-of-the-art performance on vulnerability detection benchmarks. However, due to their immense computational requirements, they are predominantly accessed via APIs or require extensive high-performance computing clusters for local deployment.
Figure 5 visualizes the relationship between model scale and adaptation techniques, e.g., fine-tuning, parameter-efficient fine-tuning, and prompt engineering; highlighting that tiny and small models are often fine-tuned in a full-parameter setting while large models are mainly prompted.
4.3.2. Hybrid (S2)